Best Plastics produces a line of cell phone covers. Its best-selling cover is the Cobra model, a slim but very durable blackiv) In columns D and E, we want to produce the profit values using Excels data table option. What is the exact entry in E2?

Please answer at least iv, v, and vi
Answer #1

Answer :

(a). Algebraic equation for TMC(q) and TR(q) are

TMC(Q) = Annual fixed cost + Variable cost *q

= 234000+2q

= 3.5q

(b). TP(q) - TMC(q)

= 3.5q - (234000 + 2q)

= 1.5q - 234000

(c). Cell entries are following

i)

B13 =B7*B10

ii)

B14 =B4+B5*B10

iii)

B15 =B13-B14

iv)

Keep E2 blank , because it's a one-way data table

v)

Entry in "Row input cell" is nothing. keep it blank

vi)

Entry in "Column input cell" is $B$10

it refers to the quantity

d)

Break-even quantity = Annual fixed cost / (selling price - variable cost)

= 234000 / (3.5-2)

= 156,000 units

e)

i)

entry in "Set cell" is C15

ii)

entry in "To value" is 0

iii)

entry in "By changing cell" is D4

Please Up-vote Thank You !!!